### v1

# Power $\Sigma_{1}$ in Card with two Woodin cardinals

Abstract
Väänänen and Welch asked in the paper [3] When cardinals determine the power set: inner models and Härtig quantifier logic which large cardinals are consistent with the power set operation $x\mapsto\mathcal{P}(x)$ being $\Sigma_{1}$-definable in the predicate $\mathrm{Card}$ of all cardinals. We show that, relative to large cardinals, this property is consistent with the existence of two Woodin cardinals.
In [3], Jouko Väänänen and Philip Welch examine the relationship between large cardinals and the $\Sigma_{1}^{\mathrm{Card}}$-definability of the power set operation, especially up to one Woodin cardinal. Here one says that power is $\Sigma_{1}^{\mathrm{Card}}$ iff there is a $\Sigma_{1}$ formula $\varphi$ of the language of set theory such that for all $x,y$, we have
$y=\mathcal{P}(x)=\{A\bigm{|}A\subseteq x\}$
iff
$\exists\alpha\in\mathrm{OR}\ \big{[}\varphi(x,y,\mathrm{Card}\cap\alpha)\big{]}.$
They show (it is an easy corollary of [3, Lemma 3.7]) that it is consistent relative to large cardinals that a Woodin cardinal exists and power is $\Sigma_{1}^{\mathrm{Card}}$. In fact, they show that if $M_{1}^{\#}$ exists and is $(0,\omega_{1}+1)$-iterable, then $M_{1}\models$ “power is $\Sigma_{1}^{\mathrm{Card}}$”. Recall here that $M_{1}$ is the canonical minimal proper class mouse with a Woodin cardinal.
They go on to ask in [3, Question 4.5] about which large cardinals are consistent with power being $\Sigma_{1}^{\mathrm{Card}}$. We make a step forward here as follows:
0.1 Theorem.
Suppose that $M_{2}^{\#}$ exists and is $(0,\omega_{1}+1)$-iterable. Then there is a proper class transitive inner model which models ZFC + “There are 2 Woodin cardinals” + “power is $\Sigma_{1}^{\mathrm{Card}}$”.
0.2 Remark.
Note that the conclusion of the theorem is not simply that $M_{2}$ satisfies the statement; our proof will not give this.
Proof of Theorem 0.1.
We will use a somewat different form of argument than the one used by [3], using instead methods as in the proof of [2, Theorem 5.1], which avoid dealing (explicitly) with the core model.
The most obvious approach to adapt the arguments for $M_{1}$ to $M_{2}$ is to formulate iterability for (enough) proper segments of $M_{2}|\delta_{0}^{M_{2}}$ in a $\Sigma_{1}^{\mathrm{Card}}$ fashion (inside $M_{2}$). But it seems not immediately clear how one might do this. Let us consider formulating iterability for a $2$-small $\omega$-sound premouse $P$ such that $\rho_{\omega}^{P}=\omega$; this is anyway the most subtle case. We would like to identify in a simple manner an appropriate tree $T$ searching for a proof that $P$ is not iterable, and then our certificate of iterability would be a transitive model which can rank $T$, proving that $T$ is wellfounded. Roughly, $T$ should build an $\omega$-maximal iteration tree $\mathcal{T}$ on $P$, embedding the tree order of $\mathcal{T}$ into the ordinals, and should build witnesses to the iterability of the Q-structures used to guide $\mathcal{T}$ at limit stages, and should build a proof that $\mathcal{T}$ reaches a stage from which it cannot be continued in such a manner, or exhibit that the last model of $\mathcal{T}$ is illfounded, if it has one. The main issue here is the iterability of the Q-structures. One might try to embed these Q-structures into something which is built by $L[\mathbb{E}]$-construction by some model of the form $M_{1}^{\#}(X)$ for some large enough (but rather arbitrary) transitive set $X$. But if we do not know that $\mathbb{R}\subseteq X$, then it is not so clear that one should be able to find such an $L[\mathbb{E}]$-construction in $M_{1}^{\#}(X)$, and also, there is the challenge of first identifying the extender sequence of $M_{1}^{\#}(X)$.
A variant of this approach is to try to make use of the Martin-Solovay tree $T_{2}$ for $\Pi^{1}_{2}$. If $\{T_{2}\}$ is $\Sigma_{1}^{\mathrm{Card}}$, then we can use it to search for the Q-structures, expecting that $\Pi^{1}_{2}$-iterability identifies them. This would certainly allow us to specify a “thick enough” tree $T$. One further detail, though, is that if $P$ is not iterable, it might in general be that there is some limit length tree $\mathcal{T}$ on $P$, such that $M_{1}(M(\mathcal{T}))$ has no segment which is a Q-structure for $M(\mathcal{T})$, and so satisfies “$\delta(\mathcal{T})$ is Woodin”. But equivalently, there is some $\Pi^{1}_{2}$-iterable premouse $N$ such that $M(\mathcal{T})\triangleleft N$ and $N$ is non-$2$-small. Moreover, this occurring is proof in itself that $P$ is not iterable (since it is $2$-small, and assuming that the earlier Q-structures were all iterable). So our tree $T$ can simply produce this situation as a witness to non-iterability.
So a key question is whether $\{T_{2}\}$ is in fact $\Sigma_{1}^{\mathrm{Card}}$. The author does not know whether one can simply prove that this holds in $M_{2}$. But we can obtain it in some $M_{2}$-like inner model, and that is where we will get the power set function being $\Sigma_{1}^{\mathrm{Card}}$.
Let us fix a finite sub-theory $\varphi_{0}$ of the first order theory of $M_{2}|\delta_{1}^{+M_{2}}$, including “I am the universe of a $2$-small premouse with 2 Woodin cardinals”, plus maybe some more statements which we will add as needed later. Say that a proper class premouse $N$ is $M_{2}$-like iff $N|\delta_{1}^{+N}\models\varphi_{0}$.
Let $(\omega_{1}^{0},u_{2}^{0},\delta_{0}^{0},\delta_{0}^{+0},\delta_{1}^{0},\delta_{1}^{+0})$ be the lexicographically least tuple $(w,u,\delta,\delta^{\prime},\varepsilon,\varepsilon^{\prime})$ such that there is an $M_{2}$-like (hence proper class) premouse $N$ such that
$(w,u,\delta,\delta^{\prime},\varepsilon,\varepsilon^{\prime})=(\omega_{1}^{N},u_{2}^{N},\delta_{0}^{N},\delta_{0}^{+N},\delta_{1}^{N},\delta_{1}^{+N}).$
(This quantification over proper classes is equivalent to one over sets, since $M_{2}$-like premice have only a set-sized extender sequence.)
Let $N$ be $M_{2}$-like, witnessing the choice of $(\omega_{1}^{0},u_{2}^{0},\delta_{0}^{0},\delta_{0}^{+0},\delta_{1}^{0},\delta_{1}^{+0})$. We will prove that $N\models$ “The power set function is $\Sigma_{1}^{\mathrm{Card}}$” (given that $\varphi_{0}$ already includes the finitely many clauses we might retroactively add later), which suffices to prove the theorem.
Work from now on in $N$. (Of course, $N\models$ “I am not fully iterable”, but by our (later) choice of $\varphi_{0}$, $N$ satisfies various first-order consequences of iterability.)
For a real $x$ and an ordinal $\alpha$, let $\iota_{0}(x,{\alpha})$ denote the least $x$-indiscernible $\mu$ such that $\mu>\alpha$.
Claim 1.
Let $M$ be a set-sized $2$-small premouse such that for some $(w,u,\delta,\delta^{\prime},\varepsilon,\varepsilon^{\prime})\in(\mathrm{OR}^{M})^{<\omega}$, $M\models$“there are 2 Woodins” and
$(w,u,\delta,\delta^{\prime},\varepsilon,\varepsilon^{\prime})=(\omega_{1}^{M},u_{2}^{M},\delta_{0}^{M},\delta_{0}^{+M},\delta_{1}^{M},\delta_{1}^{+M})$
(where $\delta_{0}^{M}<\delta_{1}^{M}$ are the 2 Woodins of $M$) and $\delta_{1}^{+M}\in\mathrm{Card}$ and $M|\delta_{1}^{+M}\models\varphi_{0}$. Then $u_{2}^{M}=u_{2}$. Therefore $\{u_{2}\}$ is $\Sigma_{1}^{\mathrm{Card}}$.
Proof.
Since $M$ is $2$-small, $M=L_{\alpha}[M|\delta_{1}^{M}]$ for some $\alpha$. Moreover, since $\delta_{1}^{+M}\in\mathrm{Card}$, note that $L[M]\models$$\delta_{0}^{M},\delta_{1}^{M}$ are both Woodin”, and in fact, $M|\delta_{1}^{+M}$ is a cardinal segment of $L[M]$, and has universe $(\mathcal{H}_{\delta_{1}^{+M}})^{L[M]}$. So since $M|\delta_{1}^{+M}\models\varphi_{0}$, $L[M]$ is $M_{2}$-like.
Now $\omega_{1}^{M}=\omega_{1}$. For certainly $\omega_{1}^{M}\leq\omega_{1}$, as otherwise $M$ has a surjection $f:\omega\to\omega_{1}$. But then if $\omega_{1}^{M}<\omega_{1}$, the properties of $M$ contradict the minimality of $\omega_{1}^{0}$.
Now let us see that $u_{2}^{M}=u_{2}$. Certainly $u_{2}^{M}\leq u_{2}$, since $M$ is correct about sharps for reals, and $M\models$$u_{2}^{M}=\sup_{x\in\mathbb{R}}\iota_{0}(x,\omega_{1})$”. But like for $\omega_{1}^{M}$, if we had $u_{2}^{M}<u_{2}$, then the properties of $M$ would contradict the lexicographic minimality of $(\omega_{1}^{0},u_{2}^{0})$. ∎
Claim 2.
For any $M$ as in Claim 1, we have $T_{2}^{M}=T_{2}$. Therefore $\{T_{2}\}$ is $\Sigma_{1}^{\mathrm{Card}}$.
Proof.
We already know $u_{2}^{M}=u_{2}$ and $u_{1}^{M}=\omega_{1}^{M}=\omega_{1}=u_{1}$. Standard calculations with indiscernibles give then that $u_{n}^{M}=u_{n}$ for all $n\in[1,\omega)$. (That is, for each $n$, $u_{n+1}=\sup_{x\in\mathbb{R}}\iota_{0}(x,u_{n})$. And $M$ satisfies that the same characterization holds. But
$\{\iota_{0}(x,\omega_{1})\bigm{|}x\in\mathbb{R}\cap M\}\text{ is cofinal in }\{\iota_{0}(x,\omega_{1})\bigm{|}x\in\mathbb{R}\},$
since $u_{2}^{M}=u_{2}$ is the supremum of both of these sets. And for all reals $x,y$,
$\iota_{0}(x,\omega_{1})<\iota_{0}(y,\omega_{1})\iff\iota_{0}(x,u_{n})<\iota_{0}(y,u_{n}),$
by indiscernibility. It follows that
$\{\iota_{0}(x,u_{n})\bigm{|}x\in\mathbb{R}\cap M\}\text{ is cofinal in }\{\iota_{0}(x,u_{n})\bigm{|}x\in\mathbb{R}\},$
and so $u_{n+1}^{M}=u_{n}=$ the common supremum of these sets.)
Now since every ordinal $<u_{\omega}$ is of the form
$t^{L[x]}(x,u_{1},\ldots,u_{n})=t^{L[x]}(x,u_{1}^{M},\ldots,u_{n}^{M})$
for some $x\in\mathbb{R}\cap M$ and $n<\omega$, it follows that $T_{2}^{M}=T_{2}$. ∎
Using $T_{2}$, we can of course define the standard tree projecting to a universal $\Sigma^{1}_{3}$ set.
Let $P$ be a sound $2$-small premouse such that $\rho_{\omega}^{P}=\omega$. We define the tree $T_{P}$ as follows; it is a tree of attempts to prove that $P$ is not $(\omega,\omega_{1})$-iterable. The tree $T_{P}$ is the natural tree building a real which codes the following things:
  1. 1.
    a countable $\omega$-maximal tree $\mathcal{T}$ on $P$, together with an order-preserving embedding of $\sigma:\mathrm{lh}(\mathcal{T})\to\omega_{1}$,
  2. 2.
    for each limit $\eta<\mathrm{lh}(\mathcal{T})$, letting $Q=Q(\mathcal{T}\!\upharpoonright\!\eta,[0,\eta)^{\mathcal{T}})$, a witness $w_{Q}$ that $Q$ is above-$\eta$, $k$-$\Pi^{1}_{2}$-iterable, where $k$ is least such that $\rho_{k+1}^{Q}\leq\delta(\mathcal{T}\!\upharpoonright\!\eta)$, and $w_{Q}$ is the appropriate kind of branch through $T_{2}$,
  3. 3.
    if $\mathcal{T}$ has successor length, an infinite descending sequence through $\mathrm{OR}^{M^{\mathcal{T}}_{\infty}}$,
  4. 4.
    if $\mathcal{T}$ has limit length, a premouse $Q_{\infty}$ such that $M(\mathcal{T})\trianglelefteq Q_{\infty}$ and $Q_{\infty}\models$$\delta(\mathcal{T})$ is Woodin”, and a witness $w^{Q_{\infty}}$ showing that $Q_{\infty}$ is above-$\delta(\mathcal{T})$, $k$-$\Pi^{1}_{2}$-iterable, where either:
    1. (a)
      $Q_{\infty}$ is non-$1$-small above $\delta(\mathcal{T})$ and $k=0$, or
    2. (b)
      $Q_{\infty}$ is $1$-small above $\delta(\mathcal{T})$ and $Q_{\infty}$ is $\delta(\mathcal{T})$-sound and $k$ is least such that $\rho_{k+1}^{Q_{\infty}}\leq\delta(\mathcal{T})$
  5. 5.
    if $\mathcal{T}$ has limit length, a structure $R$ for LST such that $\mathcal{T},Q_{\infty}\in\mathrm{HC}^{R}$ and $R\models\mathsf{ZF}^{-}$ and $R\models$ “there is no $\mathcal{T}$-cofinal branch $b$ such that $Q_{\infty}\trianglelefteq M^{\mathcal{T}}_{b}$”, together with an order-preserving embedding $\pi:\mathrm{OR}^{R}\to\omega_{1}$.
Claim 3.
$P$ is $(\omega,\omega_{1})$-iterable iff $T_{P}$ is wellfounded.
Proof.
Suppose $P$ is $(\omega,\omega_{1})$-iterable, but $x$ is a branch through $T_{P}$, giving objects $\mathcal{T},\sigma$, and if $\mathcal{T}$ has limit length, $Q_{\infty},R,\pi$. Note first that $\mathcal{T}$ is formed according to the unique strategy for $P$. For given a limit $\eta<\mathrm{lh}(\mathcal{T})$, we know that $Q=Q(\mathcal{T}\!\upharpoonright\!\eta,[0,\eta)^{\mathcal{T}})$ is $k$-$\Pi^{1}_{2}$-iterable above $\delta(\mathcal{T}\!\upharpoonright\!\eta)$, for the relevant $k$. But because $P$ is $2$-small and fully iterable, this implies that $Q$ is the correct (fully iterable) Q-structure, and $[0,\eta)^{\mathcal{T}}$ is the correct branch. So if $\mathcal{T}$ had successor length, then $M^{\mathcal{T}}_{\infty}$ would be wellfounded, a contradiction. So it has limit length. But then again, $Q_{\infty}$ is $1$-small and it is the correct Q-structure, so there is a $\mathcal{T}$-cofinal branch with $Q_{\infty}\trianglelefteq M^{\mathcal{T}}_{b}$, a contradiction. (It can’t be that $Q_{\infty}$ is non-$1$-small, since if it is, then $M_{1}(M(\mathcal{T}))|\omega_{1}^{M_{1}(M(\mathcal{T}))}\trianglelefteq Q_{\infty}$, and so $M_{1}(M(\mathcal{T}))\models$$\delta(\mathcal{T})$ is Woodin”, but this contradicts the $2$-smallness of $P$.)
Conversely, suppose $P$ is not $(\omega,\omega_{1})$-iterable. Then the putative strategy for $P$, guided by Q-structures which are iterable-above-$\delta(\mathcal{T})$, cannot succeed. Taking a (countable length) tree witnessing this of minimal length, we can use it to build a branch through $T_{P}$, as desired. ∎
We may assume:
Claim 4.
$P$ is $(\omega,\omega_{1}+1)$-iterable iff $P$ is $(\omega,\omega_{1})$-iterable.
Proof.
The corresponding statement holds in $M_{2}$, so we can build in into “$M_{2}$-like”. ∎
Claim 5.
$\{N|\omega_{1}^{N}\}$ is $\Sigma_{1}^{\mathrm{Card}}$.
Proof.
By Claims 3 and 4, $N|\omega_{1}^{N}$ is the unique premouse $M$ which satisfies $\mathsf{ZFC}^{-}$ + “Every set is countable”, with $\mathrm{OR}^{M}=\omega_{1}$, and such that there is a transitive set $R\models\mathsf{ZFC}^{-}$ with $T_{2},M\in R$ and for every $P\triangleleft M$ with $\rho_{\omega}^{P}=\omega$, we have $T_{P}\in R$ and $R\models$$T_{P}$ is wellfounded”. By Claim 2, this gives a $\Sigma_{1}^{\mathrm{Card}}$ definition of $\{N|\omega_{1}^{N}\}$. ∎
Claim 6.
Let $\theta$ be a cardinal with $\omega_{1}\leq\theta<\delta_{0}^{N}$. Let $P$ be a sound premouse such that $N|\theta\triangleleft P$ and $\rho_{\omega}^{P}=\theta$. Suppose that $P$ is above-$\theta$, $(\omega,\theta^{+}+1)$-iterable. Then $P\triangleleft N$.
Proof.
Suppose not and consider the comparison of $N$ versus $P$, producing trees $\mathcal{T}$ on $N$ and $\mathcal{U}$ on $P$, stopping at stage $\theta^{+}+1$, or at the first $\alpha$ such that $E^{\mathcal{U}}_{\alpha}\neq\emptyset$ and $\mathrm{cr}(E^{\mathcal{U}}_{\alpha})<\theta$. (Our hypothesis on the iterability of $P$ ensures that this makes sense, since $N$ is sufficently self-iterable.)
Suppose that the comparison does not reach a successful termination. It can’t be that $\alpha<\theta^{+}$ (with $\mathrm{cr}(E^{\mathcal{U}}_{\alpha})<\theta$). For if this holds, then since we are using the correct strategy for $N$, and $\mathrm{lh}(E^{\mathcal{U}}_{\alpha})$ is the index of the least disagreement between $M^{\mathcal{T}}_{\alpha}$ and $M^{\mathcal{U}}_{\alpha}$, [1, Theorem 3.5] shows that $E^{\mathcal{U}}_{\alpha}\in\mathbb{E}_{+}(M^{\mathcal{T}}_{\alpha})$, contradicting that it indexes the least disagreement.
So $\alpha=\theta^{+}$. Since $\operatorname{card}(P)<\theta^{+}$, it follows that $\mathcal{U}$ uses $\theta^{+}$-many extenders, and so $\theta^{+}$ is a limit cardinal of $M^{\mathcal{U}}_{\theta^{+}}$. But $\theta$ is the largest cardinal of $N|\theta^{+}$, and it follows that $\mathcal{T}$ also uses $\theta^{+}$-many extenders. But then we get the usual kind of contradiction to termination of comparison of weasels (in this case of height $\theta^{+}$).
So the comparison must terminate successfully prior at some stage $\alpha<\theta^{+}$. Suppose $\mathcal{U}$ is non-trivial. Then (since $\mathcal{U}$ is above $\theta$) $b^{\mathcal{U}}$ drops, $b^{\mathcal{T}}$ does not, and $M^{\mathcal{T}}_{\infty}\trianglelefteq M^{\mathcal{U}}_{\infty}$. But this contradicts the fact that $\mathrm{OR}^{P}<\theta^{+}<\mathrm{OR}^{N}$. So $\mathcal{U}$ is trivial, and note then that $P\triangleleft N$, as desired. ∎
Fix $\theta,P$ as in Claim 6, but without assuming the iterability, and suppose that $\theta<\delta_{0}^{N}$. We define a tree $T_{P}$; much as before, it is a tree of attempts to prove that $P$ is not above-$\theta$ iterable. Define this $T_{P}$ just as before, except that instead of building $\mathcal{T}$ on $P$, $T_{P}$ also searches for a countable structure $\bar{P}$ and an elementary $\varrho:\bar{P}\to P$, with $\theta\in\mathrm{rg}(\varrho)$, and builds $\mathcal{T}$ on $\bar{P}$, above $\varrho^{-1}(\theta)$.
Claim 7.
Let $\theta,P$ be as above (with $\theta<\delta_{0}^{N}$). Then $P$ is above-$\theta$, $(\omega,\theta^{+}+1)$-iterable iff $T_{P}$ is wellfounded.
Proof.
If $P$ is iterable above $\theta$, then any $\bar{P}$ (as built by $T_{P}$) is also iterable above $\bar{\theta}=\varrho^{-1}(\theta)$, which implies that $T_{P}$ is wellfounded, much as before. Conversely, suppose $P$ is not iterable above $\theta$. Then we can fix a tree witnessing this, together with the canonical Q-structures, since $N|\delta_{0}^{N}$ is sufficiently closed to compute these (and $\theta<\delta_{0}^{N}$). But then we can take a countable hull of these objects to produce a branch through $T_{P}$. ∎
[…]
Claim 8.
Let $\theta,P$ be as above. Then:
  1. 1.
    If $\delta_{0}^{N}\leq\theta<\delta_{1}^{N}$ then $P$ is above-$\theta$, $(\omega,\theta^{+}+1)$-iterable iff $T^{\prime}_{P}$ is wellfounded.
  2. 2.
    If $\omega_{1}\leq\theta<\delta_{1}^{N}$ and $T^{\prime}_{P}$ is wellfounded then $P$ is above-$\theta$, $(\omega,\theta^{+}+1)$-iterable.
Proof.
This is a simplification of the version above. ∎
Claim 9.
Let $M$ be a $2$-small premouse satisfying “there are 2 Woodin cardinals and $\delta_{1}^{+}$ exists”, $M|\delta_{1}^{+M}\models\varphi_{0}$, $\delta_{1}^{+M}\in\mathrm{Card}$, and $\mathrm{Card}\cap\delta_{1}^{+M}=\mathrm{Card}^{M}\cap\delta_{1}^{+M}$, and for every cardinal $\theta<\delta_{1}^{M}$ and $P$ such that $P\trianglelefteq M$ and $\rho_{\omega}^{P}=\theta$, we have:
  1. 1.
    if $\theta<\delta_{0}^{M}$ then $T_{P}$ is wellfounded, and
  2. 2.
    if $\delta_{0}^{M}\leq\theta<\delta_{1}^{M}$ then $T^{\prime}_{P}$ is wellfounded.
Then $M|\delta_{1}^{M}=N|\delta_{1}^{N}$.
Proof.
As in the proof of Claim 1, $M$ is $M_{2}$-like, $\omega_{1}^{M}=\omega_{1}$ and $u_{2}^{M}=u_{2}$. So by minimality, $\delta_{0}^{N}\leq\delta_{0}^{M}$, and if $\delta_{0}^{N}=\delta_{0}^{M}$ (and hence $\delta_{0}^{+N}=\delta_{0}^{+M}$, by the correctness of $M$’s cardinals), then $\delta_{1}^{N}\leq\delta_{1}^{M}$, and if $\delta_{1}^{N}=\delta_{1}^{M}$ then $\delta_{1}^{+N}=\delta_{1}^{+M}$.
Now $N|\delta_{0}^{N}=M|\delta_{0}^{N}$. For let $\theta<\delta_{0}^{N}$ be a cardinal and suppose $N|\theta=M|\theta$, and let $P\triangleleft M$ with $\rho_{\omega}^{P}=\theta$. Then by hypothesis (and since $\delta_{0}^{N}\leq\delta_{0}^{M}$), $T_{P}$ is wellfounded, so by Claims 6 and 7, $P\triangleleft N$, as desired.
Now since $N|\delta_{0}^{N}=M|\delta_{0}^{M}$ and $M\subseteq N$, it follows that $\delta_{0}^{M}$ is Woodin in $M$. So $\delta_{0}^{N}=\delta_{0}^{M}$, and so $\delta_{0}^{+N}=\delta_{0}^{+M}$.
A similar calculation as before now gives that $N|\delta_{1}^{N}=M|\delta_{1}^{N}$, and then again since $M\subseteq N$, it follows that $\delta_{1}^{N}=\delta_{1}^{M}$ and $\delta_{1}^{+N}=\delta_{1}^{+M}$, as desired. ∎
Claim 10.
$\{N|\delta_{1}^{N}\}$ is $\Sigma_{1}^{\mathrm{Card}}$.
Proof.
This follows easily from the preceding claims. Our $\Sigma_{1}^{\mathrm{Card}}$ formula $\varphi(M)$ (in variable $M$) just demands that $M$ satisfies the hypothesis of Claim 9, by asserting that there is a transitive model which ranks the relevant trees $T_{P}$ and $T^{\prime}_{P}$; by earlier claims, we can identify these trees in a $\Sigma_{1}^{\mathrm{Card}}$ fashion. We then get that $\varphi(M)$ holds iff $M=N|\delta_{1}^{N}$. ∎
It now follows that $N\models$ “The power set function is $\Sigma_{1}^{\mathrm{Card}}$”, like in the proof for $L$: Given $x,y$, we have $y=\mathcal{P}(x)$ iff there is a cardinal $\theta$ such that $x,y\in L_{\theta}(N|\delta_{1}^{N})$ and $y=\mathcal{P}(x)\cap L_{\theta}(N|\delta_{1}^{N})$. This completes the proof. ∎
0.3 Question.
Suppose $M_{2}^{\#}$ exists and is $(0,\omega_{1}+1)$-iterable. Does $M_{2}$ satisfy “power is $\Sigma_{1}^{\mathrm{Card}}$?
0.4 Question.
Is the theory ZFC + “there are 2 Woodin cardinals and a measurable above” + “power is $\Sigma_{1}^{\mathrm{Card}}$” consistent (relative to large cardinals)?
Acknowledgements
This research was funded by the Austrian Science Fund (FWF) [10.55776/Y1498].